Chicago Cola is considering the purchase of a special-purpose bottling machine for $24,000. It is expected to have a useful l2. Payback period. (Round your answer to two decimal places.) The payback period is 2.85 years. 3. Discounted payback period.

O Data Table - X Year $ Amount 11,000 9,000 7,000 4,000 $ 31,000 Total Print Done
Answer #1

Solution

Year Annual Cost Savings CFAT Cumulative CFAT PVIF @ 18% Discounted CFAT Cumulative Discounted CFAT
1 $                   11,000.00 $ 11,000.00 $            11,000.00 0.8475 $              9,322.03 $                                        9,322.03
2 $                     9,000.00 $   9,000.00 $            20,000.00 0.7182 $              6,463.66 $                                      15,785.69
3 $                     7,000.00 $   7,000.00 $            27,000.00 0.6086 $              4,260.42 $                                      20,046.11
4 $                     4,000.00 $   4,000.00 $            31,000.00 0.5158 $              2,063.16 $                                      22,109.27
Total Discounted CFAT $            22,109.27
Less: Initial Investment $          -24,000.00
NET PRESENT VALUE $             -1,890.73

(1) Net Present Value will be $ (-) 1,890.73

(2) Cash Payback Period: From the Cumulative CFAT column, it can be observed that Initial Outflow of $ 24,000 will be recovered in 2 to 3 years. So, assuming at Xth year the Cumulative Cash Flow should be $ 24,000, value of X using Interpolation,

(X-2)/(3-2) = (24000-20000)/(27000-20000)

Or, (X-2) = 4000/7000

Or, X-2 = 0.57 (Approx)

Or, X = 2.57

Therefore Cash Payback Period = 2.57 Years

(3) Discounted Payback Period: If referred to the Cumulative Discounted CFAT column, it is clearly seen that the company is not able to recover its initial outlay in the given time frame. Hence, Discounted Payback Period could not be calculated.

Note: The CFAT is actually Annual Cost Savings, not revenue. Hence the depreciation effect has nothing to do with this calculation.